OpenStudy (anonymous):

How do we use the zero property to solve quadratic equations? Please give me an example

OpenStudy (anonymous):

you may need to be more specific.

OpenStudy (anonymous):

I dont know how more specific I can be...? I'm basically asking how do we use zero property to solve a quadratic equation and I just want a example.

OpenStudy (anonymous):

kk lets see, the zero property says if ab=0 then a=0 or b=0 or they both equal 0. Consider this \[x ^{2}=5x\]

OpenStudy (anonymous):

then x^2-5x=0, factorise x(x-5)=0, so by the zero property x=0 or(x-5)=0 in which case x=5. Hence x=0 or x=5

OpenStudy (anonymous):

Alright great example thanks for helping me

OpenStudy (anonymous):

no problemo